# Frustration-free free fermions

### v1

@@ Abstract
Notice that $e_{\bm{k}}=0$ at the K and K’ points ${\bm{k}}=(\pm 1/3,\pm 1/3)$. This implies the sets of localized orbitals $\{\hat{\psi}^{\dagger}_{\bm{R}}\}$ and $\{\hat{\phi}^{\dagger}_{\bm{R}}\}$ are not linearly independent over the full lattice in general, and correspondingly ${\bm{\psi}}_{{\bm{k}}}$ and ${\bm{\phi}}_{{\bm{k}}}$ are undefined at ${\bm{k}}$ = K and K’.
In momentum space, one finds ${\hat{H}=\sum_{\bm{k}}\hat{\bm{c}}^{\dagger}_{{\bm{k}}}h({\bm{k}})\hat{\bm{c}}_{{\bm{k}}}}+L^{2}$, where the Bloch Hamiltonian reads $h({\bm{k}})={t({\bm{\psi}}_{{\bm{k}}}{\bm{\psi}}_{{\bm{k}}}^{\dagger}-{\bm{\phi}}_{{\bm{k}}}{\bm{\phi}}_{{\bm{k}}}^{\dagger})}$. As condition Eq. (5) implies ${\bm{\psi}}_{{\bm{k}}}^{\dagger}{\bm{\phi}}_{{\bm{k}}}=0$, for $e_{\bm{k}}\neq 0$ the normalized eigenvectors for the conduction and valence bands are respectively $\bm{u}_{\bm{k},c}={\bm{\psi}}_{{\bm{k}}}/\sqrt{e_{\bm{k}}}$ and $\bm{u}_{\bm{k},v}={\bm{\phi}}_{{\bm{k}}}/\sqrt{e_{\bm{k}}}$. The associated fermion creation operators are $\hat{f}^{\dagger}_{\bm{k},\alpha}=\hat{\bm{c}}_{\bm{k}}^{\dagger}\bm{u}_{\bm{k},\alpha}$ for $\alpha=c,v$. The band dispersions ±$\pm te_{\bm{k}}$ touch quadratically at $\bm{k}=$ K and K’ (Fig. 1(c)), as required by the F4 condition. We remark that our F4 Hamiltonian can also be viewed as the consequence of tuning the Dirac velocity to zero through the introduction of a chiral-symmetric third-nearest neighbor hopping (Appendix B).
Figure 1: Frustration-free free-fermion model on the honeycomb. (a) Lattice convention and the definition of the local orbital corresponding to the conduction band. (b) Local orbital for the valence band. (c) Band structure. (d) Non-zero eigenvalue of the quantum metric tensor in arbitrary units, which diverges as $1/\delta q^{2}$ around the K point. (e) Excitation energy of density fluctuations within the single-mode approximation, which acquires an anomalous gap $O(1/\log L)$ due to the singular quantum metric. We take $\delta\bm{k}=(\bm{b}_{1}+\bm{b}_{2})/L$. Crosses are computed values and the dashed line indicates an inverse-log fit with fitting parameters reported in Appendix E.
@@
Interaction and density fluctuations— Does the same excitation gap scaling hold beyond the non-interacting limit? For a general fermionic frustration-free model, we have shown in Ref. [53] that, assuming a power-law decay for a two-point correlation function in the ground state, the Gosset-Huang inequality [56] implies $\epsilon(L)\sim O((\log L)^{2}/L^{2})$. We now turn to validating this assertion when a local repulsion $\hat{U}=U\sum_{\bm{R}}(\hat{\psi}^{\dagger}_{\bm{R}}\hat{\psi}_{\bm{R}})(\hat{\phi}^{\dagger}_{\bm{R}}\hat{\phi}_{\bm{R}})$ with $U>0$ is added to the F4 model. By design, each local term in $\hat{U}$ is positive semi-definite and annihilates the non-interacting ground state $|\Phi\rangle$, and so $\hat{H}^{\rm int}=\hat{H}+\hat{U}$ remains a frustration-free Hamiltonian. Yet, $\hat{U}$ modifies the excitation energies although the ground states are unchanged, and so the behavior of $\epsilon(L)$ should be reexamined.
@@
To bound $\epsilon(L)$, it suffices to consider suitable variational excited states. In the context of interacting frustration-free bosonic systems, this is often achieved through the density fluctuations treated within the single-mode approximation (SMA) [57]. The trial states take the form $|{\bm{k}}\rangle\propto\sum_{{\bm{R}}}\hat{n}_{\bm{R}}e^{i{\bm{k}}\cdot{\bm{R}}}|\Phi\rangle$, where $\hat{n}_{\bm{R}}$ is the total density operator in the unit cell ${\bm{R}}$. The SMA energy is given by the energy expectation value $E^{\rm SMA}({\bm{k}})=\langle{\bm{k}}|\hat{H}|{\bm{k}}\rangle/{\langle{\bm{k}}|{\bm{k}}\rangle}$, and corresponds to a quadratic gapless mode $E^{\rm SMA}({\bm{k}})\sim O(|{\bm{k}}|^{2})$ in many known models [58, 59, 60]. This quadratic dispersion again translates into a finite-size excitation gap $\epsilon(L)\sim O(1/L^{2})$, consistent with the bound obtained through the Gosset-Huang inequality [50].
@@
Curiously, we now show that this commonly used ansatz does not provide a useful bound even in the free-fermion limit. Specifically, we show that $E^{\rm SMA}$ acquires an anomalous gap of $O(1/\log L)$ due to the diverging quantum metric in the F4 model. For definiteness, we take $|\Phi\rangle$ to be the state obtained by filling all the negative-energy modes. As a technical trick, we further assume a finite linear size $L$ which is not divisible by $3$. This avoids sampling the singular K and K’ points in the finite-size momentum mesh and leads to a unique ground state  [^1] The case with $L$ divisible by $3$ can also be handled. The ground state degeneracy leads only to an $O(1)$ extra term compared to the $O(\log L)$ divergence of the denominator in the SMA energy, and so it is inessential to our discussion.. The local density operator is given by $\hat{n}_{\bm{R}}=\sum_{\sigma=A,B}\hat{c}_{{\bm{R}},\sigma}^{\dagger}\hat{c}_{{\bm{R}},\sigma}$. In the momentum-space, for $\bm{k}\neq\bm{0}$ and using $\hat{f}_{\bm{q},c}|\Phi\rangle=\hat{f}_{\bm{q},v}^{\dagger}|\Phi\rangle=0$, we have
@@
Now, we notice that $\bm{u}_{\bm{q},\alpha}$ are not well-defined at the K and K’ points at which the two bands touch. At a momentum $\delta\bm{q}$ away from these singular points, the non-zero eigenvalue of the quantum metric diverges as $\sim 1/\delta q^{2}$ (Appendix C). When integrated, $\int d^{2}(\delta\bm{q})\,\delta q^{-2}$ diverges logarithmically with the low-momentum cutoff $1/L$. This then enters into the normalization factor in Eq. (11), which goes as $\langle\delta\bm{k}|\delta\bm{k}\rangle\sim\log(L)\delta k^{2}$. Although the divergence is cured in the numerator, which goes as $O(1)\delta k^{2}$, the SMA energy acquires an anomalous gap which vanishes logarithmically in the thermodynamic limit,
$\begin{split}E^{\rm SMA}(\delta{\bm{k}})=\frac{\langle\delta\bm{k}|\hat{H}|\delta\bm{k}\rangle}{\langle\delta\bm{k}|\delta\bm{k}\rangle}\approx\frac{O(1)}{\log L}+\cdots.\end{split}$ (13)
Therefore, in contrast to many gapless frustration-free bosonic models [58, 59, 60], we cannot conclude the asserted $O(1/L^{2})$ finite-size scaling through $E^{\rm SMA}(\bm{k})$.
Figure 2: Energy of particle-hole excitations in $\hat{H}^{\rm int}$ at $U=2t$. (a) Energy spectrum of $\hat{H}^{\rm int}$ restricted into the subspace spanned by $|\bm{k},\bm{q}\rangle$ for $\bm{k}$ along the M-$\Gamma$-K line computed for $L=50$. The boundary of the particle-hole continuum can be estimated as $\min_{\bm{q}}(te_{\bm{q}}+(t+U)e_{\bm{q}+\bm{k}})$ (green solid line), which becomes exact in the non-interacting limit with $U=0$ (grey dashed line). Note that interaction $U>0$ modifies the excitation energies although it keeps the ground states unchanged. (b) System size dependence of the particle-hole energy at $\bm{k}=\bm{0}$. The dashed line shows the $O(1/L^{2})$ bound in Eq. (14).
Nevertheless, we can still demonstrate $\epsilon(L)\sim O(1/L^{2})$ using different variational excited states. In our fermionic context, another natural candidate is the zero-momentum particle-hole excitation $|\bm{0},\bm{q}\rangle$. In the free-fermion model, the smallest excitation energy is given by $\bm{q}$ in the vicinity of the K and K’ points, and it takes the form $2te_{\bm{q}_{L}}^{\rm min}$ with $e_{\bm{q}_{L}}^{\rm min}=\min\{e_{\bm{q}}>0\}\sim O(1/L^{2})$. As we show in Appendix D, a similar bound of $(2t+U)e_{\bm{q}_{L}}^{\rm min}$ can be derived when $U>0$. Analyzing $e_{\bm{q}_{L}}^{\rm min}$ further, we obtain
which provides the desired $O(1/L^{2})$ bound.
In Fig. 2, we show that the the particle-hole excitation indeed disperses quadratically around $\Gamma$, and verifies the discussed finite-size bound on the minimum excitation energy. Numerically, the bound of the charge-neutral finite-size gap using these ansatz states decay as $L^{-1.99(2)}$ (Appendix E), consistent with our analytical results.
Conclusions—In this work, we derived the necessary and sufficient conditions for local frustration-free free-fermion systems, and proved that the touching between conduction and valence bands in such model must be quadratic or softer. We also provided a concrete model on the honeycomb lattice, and found that the diverging quantum metric led to an anomalous gap in the density fluctuations. Nevertheless, one could construct charge-neutral excitations which bound the finite-size scaling of the excitation gap to be $O(1/L^{2})$ even in the presence of interaction, consistent with general results we obtained based on the ground-state correlation functions in an accompany work, Ref. [53].
Acknowledgements.
We thank Hosho Katsura, Ken Shiozaki, Hal Tasaki, Carolyn Zhang, and Ruben Verresen for useful discussions. H.C.P. acknowledges support from the National Key R&D Program of China (Grant No. 2021YFA1401500) and the Hong Kong Research Grants Council (C7037-22GF). The work of H.W. is supported by JSPS KAKENHI Grant No. JP24K00541. S.O. was supported by RIKEN Special Postdoctoral Researchers Program. H.W., S.O., and H.C.P. thank the Yukawa Institute for Theoretical Physics, Kyoto University, where we started the project during the YITP workshop YITP-T-24-03 on “Recent Developments and Challenges in Topological Phases.”
Appendix A Details of the Dirac trick
Appendix B Tight-binding form
Consider a chiral symmetric honeycomb model of spinless fermions. Let the nearest-neighbor hopping be $t_{1}$. Assuming chiral symmetry represented by $\sigma_{3}$ in the sublattice space, the next allowed hopping $t_{2}$ is between third nearest neighbors (Fig. 3). In momentum space, the Bloch Hamiltonian is given by $h_{\bm{k}}=Q_{\bm{k}}\sigma^{+}+Q_{\bm{k}}^{*}\sigma^{-}$, where
Figure 3: Spinless fermions on the honeycomb lattice. (a) The shortest bonds on the honeycomb lattice assuming a chiral symmetry forbidding intra-sublattice hopping. (b) Brillouin zone and special momenta.
Appendix C Quantum metric of the model and variational energy of density fluctuations
While the quantum metric is well-defined for isolated bands, it could become singular at band touching points. Indeed, in our F4 honeycomb model one finds $g_{ij}(\bm{q})\sim O(1/\delta q^{2})$ when $\bm{q}=\bm{q}_{\rm K}+\delta\bm{q}$ is around the K point (and similarly for the K’ point). More explicitly, we can parameterize $\bm{q}=\bm{q}_{K}+\delta q(\cos\theta,\sin\theta)$ and compute the leading singularity as
@@ Appendix C Quantum metric of the model and variational energy of density fluctuations
We can now study how this singularity affects the SMA energy. For the energy expectation value in Eq. (11), the leading term in the $\delta k\rightarrow 0$ limit is given by $t\sum_{\bm{q}}(2e_{\bm{q}})g_{ij}(\bm{q})\delta k_{i}\delta k_{j}$. Since $e_{\bm{q}}\sim O(\delta q^{2})$ around the K and K’ points, the divergence in the quantum metric is compensated and one finds $\langle\delta\bm{k}|\hat{H}|\delta\bm{k}\rangle\sim O(1)\delta k^{2}$, agreeing with general results derived form a double-commutator bound assuming only the locality and charge conservation symmetry of $\hat{H}$ [57, 58, 59, 60]. In typical SMA calculations, the normalization factor starts at $O(1)$ and so the SMA energy is gapless and disperses quadratically.
@@
[…]
@@
Therefore, although both the numerator and denominator in the SMA energy start with a quadratic term $\sim\delta k^{2}$, the coefficient for the numerator is $O(1)$ whereas that for the denominator is $O(\log L)$. This leads to $E^{\rm SMA}(\delta\bm{k})\sim O(1/\log L)$ in the limit $\delta k\rightarrow 0$, giving rise to an anomalous gap.
Appendix D Interacting bound of the neutral excitation gap
We first Fourier transform the interaction $\hat{U}$ to find
@@ Appendix D Interacting bound of the neutral excitation gap
Next, we consider how the interaction modifies the charge-neutral excitation gap. While the particle-hole excitations $|\bm{k},\bm{q}\rangle$ are eigenstates of the free-fermion Hamiltonian, they become coupled when the interaction $\hat{U}$ is added to the Hamiltonian. As translation symmetry is preserved, sectors labeled by $\bm{k}$ remain decoupled. Let $\hat{H}^{\rm ph}|_{\bm{k}=\bm{0}}$ be the restriction of the Hamiltonian into the zero-momentum sector, defined by the matrix elements $\langle\bm{0},\bm{q}|\hat{H}^{\rm int}|\bm{0},\bm{q}^{\prime}\rangle$. Fig. 2(b) plots the low-lying eigenvalues of $\hat{H}^{\rm ph}|_{\bm{k}=\bm{0}}$ for different system sizes $L$. We can now find an upper bound to the smallest eigenvalue through
$\begin{split}\min{\rm eig}\left(\hat{H}^{\rm ph}|_{\bm{k}=\bm{0}}\right)\leq&\min_{\bm{q}}\langle\bm{0},\bm{q}|\hat{H}^{\rm int}|\bm{0},\bm{q}\rangle\\ =&\min_{\bm{q}}\left(2te_{\bm{q}}+Ue_{\bm{q}}-\frac{U}{V}e_{\bm{q}}^{2}\right)\\ \leq&(2t+U)\min_{\bm{q}}e_{\bm{q}},\end{split}$ (23)

### v2

@@ Abstract
Notice that $e_{\bm{k}}=0$ at the K and K’ points ${\bm{k}}=(\pm 1/3,\pm 1/3)$ and so ${\bm{\psi}}_{{\bm{k}}}$ and ${\bm{\phi}}_{{\bm{k}}}$ vanish there. This implies that the sets of localized orbitals $\{\hat{\psi}^{\dagger}_{\bm{R}}\}$ and $\{\hat{\phi}^{\dagger}_{\bm{R}}\}$ are not linearly independent over the full lattice.
In momentum space, ${\hat{H}=\sum_{\bm{k}}\hat{\bm{c}}^{\dagger}_{{\bm{k}}}h({\bm{k}})\hat{\bm{c}}_{{\bm{k}}}}+L^{2}$ with the Bloch Hamiltonian $h({\bm{k}})={t({\bm{\psi}}_{{\bm{k}}}{\bm{\psi}}_{{\bm{k}}}^{\dagger}-{\bm{\phi}}_{{\bm{k}}}{\bm{\phi}}_{{\bm{k}}}^{\dagger})}$. As condition Eq. (5) implies ${\bm{\psi}}_{{\bm{k}}}^{\dagger}{\bm{\phi}}_{{\bm{k}}}=0$, for $e_{\bm{k}}\neq 0$ the normalized eigenvectors for the conduction and valence bands are respectively $\bm{u}_{\bm{k},c}={\bm{\psi}}_{{\bm{k}}}/\sqrt{e_{\bm{k}}}$ and $\bm{u}_{\bm{k},v}={\bm{\phi}}_{{\bm{k}}}/\sqrt{e_{\bm{k}}}$. The associated fermion creation operators are $\hat{f}^{\dagger}_{\bm{k},\alpha}=\hat{\bm{c}}_{\bm{k}}^{\dagger}\bm{u}_{\bm{k},\alpha}$ for $\alpha=c,v$. The band dispersions $\pm te_{\bm{k}}$ touch quadratically at $\bm{k}=$ K and K’ (Fig. 1(c)), as required by the F4 condition. Furthermore, taking the limit $\bm{k}\to\text{K},\text{K}^{\prime}$ reveals that the normalized Bloch states $\bm{u}_{\bm{k},\alpha}$ are ill-defined at these gap closing points, indicating that the quadratic band touching is singular [57] .
Figure 1: Frustration-free free-fermion model on the honeycomb. (a) Lattice convention and the definition of the local orbital corresponding to the conduction band. (b) Local orbital for the valence band. (c) Band structure. (d) Non-zero eigenvalue of the quantum metric tensor in arbitrary units, which diverges as $1/\delta q^{2}$ around the K point. (e) Excitation energy of density fluctuations within the single-mode approximation, which acquires an anomalous gap $O(1/\log L)$ due to the singular quantum metric. We take $\delta\bm{k}=(\bm{b}_{1}+\bm{b}_{2})/L$. Crosses are computed values and the dashed line indicates an inverse-log fit with fitting parameters reported in Appendix B.
@@
Interaction and density fluctuations— Does the same excitation gap scaling hold beyond the non-interacting limit? For a general fermionic frustration-free model, we have shown in Ref. [53] that, assuming a power-law decay of a two-point correlation function in the ground state, the Gosset-Huang inequality [58] implies $\epsilon(L)=O((\log L)^{2}/L^{2})$. We now turn to validating this assertion when a local repulsion $\hat{U}=U\sum_{\bm{R}}(\hat{\psi}^{\dagger}_{\bm{R}}\hat{\psi}_{\bm{R}})(\hat{\phi}^{\dagger}_{\bm{R}}\hat{\phi}_{\bm{R}})$ with $U>0$ is added to the F4 model. By design, each local term in $\hat{U}$ is positive semi-definite and annihilates the non-interacting ground state $|\Phi\rangle$, and so $\hat{H}^{\rm int}=\hat{H}+\hat{U}$ remains a frustration-free Hamiltonian. Yet, $\hat{U}$ modifies the excitation energies and so the behavior of $\epsilon(L)$ should be reexamined.
@@
To bound $\epsilon(L)$, it suffices to consider suitable variational excited states. In the context of interacting frustration-free bosonic systems, this is often achieved through the density fluctuations treated within the single-mode approximation (SMA) [59]. The trial states take the form $|{\bm{k}}\rangle\propto\sum_{{\bm{R}}}\hat{n}_{\bm{R}}e^{i{\bm{k}}\cdot{\bm{R}}}|\Phi\rangle$, where $\hat{n}_{\bm{R}}$ is the total density operator in the unit cell $\bm{R}$}. The SMA energy is given by the energy expectation value $E^{\rm SMA}({\bm{k}})=\langle{\bm{k}}|\hat{H}|{\bm{k}}\rangle/{\langle{\bm{k}}|{\bm{k}}\rangle}$ and satisfies $E^{\rm SMA}({\bm{k}})=O(|{\bm{k}}|^{2})$ in many known models [60, 61, 62, 50].
Surprisingly, this commonly used ansatz does not provide a useful bound in the present honeycomb model. Instead, due to a diverging quantum metric in the vicinity of the gap closing point of the non-interacting limit, $E^{\rm SMA}$ acquires an anomalous gap of $O(1/\log L)$ [53]. Nevertheless, we can still demonstrate $\epsilon(L)=O(1/L^{2})$ using variational excited states in Eq. (9). Consider zero-momentum particle-hole excitations $|\bm{0},\bm{q}\rangle$. When $U=0$, the excitation energy for $\bm{q}$ around the K and K’ points takes the form $2te_{\bm{q}_{L}}^{\rm min}$ with $e_{\bm{q}_{L}}^{\rm min}=\min\{e_{\bm{q}}>0\}=O(1/L^{2})$. As shown in Appendix B, a similar bound of $(2t+U)e_{\bm{q}_{L}}^{\rm min}$ can be derived when $U>0$. Analyzing $e_{\bm{q}_{L}}^{\rm min}$ further, we obtain
which provides the desired $O(1/L^{2})$ bound.
Figure 2: Energy of particle-hole excitations in $\hat{H}^{\rm int}$ at $U=2t$. (a) Energy spectrum of $\hat{H}^{\rm int}$ restricted to the subspace spanned by $|\bm{k},\bm{q}\rangle$ for $\bm{k}$ along the M-$\Gamma$-K line computed for $L=50$. The boundary of the particle-hole continuum can be estimated as $\min_{\bm{q}}(te_{\bm{q}}+(t+U)e_{\bm{q}+\bm{k}})$ (green solid line), which becomes exact in the non-interacting limit with $U=0$ (grey dashed line). Note that interaction $U>0$ modifies the excitation energies although it keeps the ground states unchanged. (b) System size dependence of the particle-hole energy at $\bm{k}=\bm{0}$. The dashed line shows the $O(1/L^{2})$ bound in Eq. (10).
In Fig. 2, we show that the particle-hole excitation indeed disperses quadratically around $\Gamma$, confirming the discussed finite-size bound on the minimum excitation energy. Numerically, the corresponding bound of the charge-neutral finite-size gap decays as $L^{-1.99(2)}$ (Appendix B), consistent with our analytical results.
Correlation functions— Having derived the general condition and illustrated it through a 2D honeycomb model, we next demonstrate how the F4 conditions inform the general behavior of ground state correlation functions.
Generally, the ground state correlation functions of a gapped, local Hamiltonian exhibit exponential decay [63, 64]. Conversely, while a power-law decay is naturally anticipated in the ground state of a gapless system, there can also be gapless models in which all ground state correlation functions decay exponentially with distance. Both possibilities are possible in frustration-free models. For instance, the uncle Hamiltonian  [65, 50] for matrix product states is gapless but its ground state has only exponentially decaying correlation functions. In contrast, the ground state of the Motzkin chain [66, 67] exhibits power-law correlation functions with critical exponents $z\geq 2$.
We now investigate how the F4 condition informs the behavior of the ground-state correlations. A free-fermion model is gapless if and only if there is a fermion mode with energy $\epsilon=0$ in the thermodynamic limit. While a zero-energy mode implies ground state degeneracy, one can ask: does there exist one ground state exhibiting only exponentially decaying correlation functions? We say the gapless model is non-critical when such a ground state exists.
A free-fermion model must be non-critical if there is a continuous band gap separating the conduction and the valence bands in the thermodynamic limit. It remains to consider the case when there is a band touching point at $\epsilon=0$. From our proof, we know that a ground state $|\Phi\rangle$ of any F4 Hamiltonian can be constructed by occupying all the locally defined $\phi$ orbitals. Let $\epsilon_{v}^{*}$ denote the top of the valence bands, defined as the maximum single-particle energy among the valence bands over the continuous BZ. When $\epsilon_{v}^{*}=0$, the $\phi$ orbitals are incomplete in the following sense: Suppose there are $N_{v}$ valence bands. Let $V$ be the system volume. From the general discussion on the excitation gap, we see that $\epsilon_{v}^{*}=0$ if and only if there are fewer than $VN_{v}$ linearly independent orbitals among the set $\{\hat{\phi}_{\bm{R},\beta}\}_{\bm{R},\beta}$.
Importantly, incompleteness of the $\phi$ orbitals is not equivalent to criticality: If there exist additional locally defined orbitals $\hat{\phi}^{\prime}_{\bm{R}^{\prime}\beta^{\prime}}$ satisfying $\{\hat{\phi}^{\prime}_{\bm{R}^{\prime}\beta^{\prime}},\hat{\psi}_{\bm{R}\alpha}\}=0$, they can be added to the original Hamiltonian without altering the conduction bands. When such $\phi^{\prime}$ orbitals can be found such that the combined set of $\phi$ and $\phi^{\prime}$ orbitals contains $VN_{v}$ linearly independent operators, the modified F4 Hamiltonian yields $\epsilon_{v}^{*}<0$. Since the ground state of the modified, gapped Hamiltonian is also a ground state of the original gapless model, we conclude the original model is non-critical. The desired $\phi^{\prime}$ orbitals may not exist. For instance, in the case of the nearest-neighbor kagome model, the set of flat-band orbitals can only be completed by including extended states that are delocalized along at least one direction of the system [15].
Naturally, the feasibility to complete the set of $\phi$ orbitals is tied to the analytical properties of the Bloch wave functions around the gap closing point. Recall, our F4 condition implies any gapless F4 model can be deformed to a flat-valence-band limit by setting $\nu_{{\bf R}\beta}\rightarrow 0$. This allows us to analyze the criticality problem in the framework of flat-band singular band touching [57]. When there is a singular band touching point, the Bloch states in its vicinity must be non-analytic the ground state correlation functions exhibit power-law decay [53]. Conversely, for a single-valence-band problem with only non-singular band touching points, Ref. [57] showed that a band gap can open without modifying the Bloch states. Such scenario corresponds to non-critical gapless F4 model, which is generally the case for one-dimensional models [53].
The analysis above motivates the following conjecture: a gapless F4 model is non-critical if and only if, in the flat-band limit, all the band touchings between the conduction and valence bands are non-singular. While the “only if” direction has been established [53], the “if” part is currently known only for the special case of a single valence band [57].
Gapped topological bands— We now turn to gapped F4 models. It is well-known that the Kitaev model for a 1D topological superconductor admits a F4 realization. What other forms of band topology are compatible with the F4 condition? For concreteness, we assume there is a band gap and the bottom of the conduction bands, defined in the thermodynamic limit, satisfies $\epsilon_{c}^{*}>0$. We now show that the only possible tenfold-way topological phases realizable in the valence bands would be weak phases obtained by stack of 1D phases. We further conjecture that any other stable crystalline topological phases are similarly forbidden. In addition, we show that fragile topology is generally compatible with the F4 conditions.
We argue by first addressing the possible topology of the conduction bands. Given the gap condition $\epsilon_{c}^{*}>0$, the $\psi$ orbitals in the defining F4 Hamiltonian provide a set of compact orbitals which span all the states in the conduction band. Note that these orbitals might be over-complete, which happens when the number of $\psi$ orbitals in each unit cell exceeds the number of conduction bands. Nevertheless, Ref. [68] proved that the only ten-fold way topology realizable in such bands are weak phases obtained by stacking 1D topological phases. As the stable topological indices for the whole system is necessarily zero, we conclude the only possible ten-fold way topology of the valence bands will also be weak phases stemming from the 1D classification. Although not addressed explicitly in Ref. [68], the analysis therein could be enriched by additional crystalline symmetries, and so we conjecture that other forms of stable topology beyond those deformable to representatives built from 0D and 1D topological building blocks [69, 70, 71, 72, 73, 74, 75, 76] are similarly forbidden in gapped F4 Hamiltonians with a gap $\epsilon_{c}^{*}>0$.
Fragile topology [77], which arises frequently as the complement of atomic bands, could still exist even when stable topology is ruled out. Indeed, when one takes the flat-band limit for the fragile topological model in Ref. [78] for twisted bilayer graphene, the resulting Hamiltonian satisfies our F4 condition. In Ref. [53], we further demonstrate that the valence bands could acquire dispersion without breaking the F4 conditions.
Conclusions and outlook—In this work, we derived the necessary and sufficient conditions for local frustration-free free-fermion systems, and proved that the touching between conduction and valence bands in such models must be quadratic or softer. We also provided a concrete model on the honeycomb lattice, and found that the diverging quantum metric led to an anomalous gap in the density fluctuations. Nevertheless, one could construct charge-neutral excitations which bound the finite-size scaling of the excitation gap to be $O(1/L^{2})$ even in the presence of interactions, consistent with the general results in Ref. [53]. We further show that gapless F4 models can be non-critical, and that the criticality is tied to the presence of singular band touching points in the flat-band limit [57]. In addition, under a gap condition, we show that the valence bands of a F4 model can only be atomic, weak-topological with 1D building blocks, or fragile topological.
Our results also lead to a collection of no-go concerning the possibility of realizing various topological phases using F4 Hamiltonains. The dispersion bound rules phases like Dirac and Weyl semimetals with linear dispersion. Under the gap condition  [^1] We remark that our conditions leave open one technical possibility: if a continuous band gap exists between the conduction and valence bands although $\epsilon_{v}^{*}=\epsilon_{c}^{*}=0$, and that both sets cannot be made complete by supplying symmetric, local orbitals, then our analysis is silent on the band topology which could be realized. , we also rule out the realization of any Wannier-obstructing tenfold-way topology [80, 81, 82], such as Chern and $\mathbb{Z}_{2}$-topological insulators [83, 84, 85, 68]. However, we remark that once interactions are included, frustration-free fermionic models can host nontrivial topology forbidden in the free-fermion limit. This has been demonstrated by the construction of commuting projector Hamiltonians for the $\mathbb{Z}_{2}$ topological insulator [86] and recent developments in fermionic projected entangled pair states (PEPS) [87, 88].
Acknowledgements.
We thank Hosho Katsura, Ken Shiozaki, Hal Tasaki, Carolyn Zhang, Ruben Verresen, and Shenghan Jiang for useful discussions. H.C.P. acknowledges support from the National Key R&D Program of China (Grant No. 2021YFA1401500) and the Hong Kong Research Grants Council (C7037-22GF). The work of H.W. is supported by JSPS KAKENHI Grant No. JP24K00541. S.O. was supported by RIKEN Special Postdoctoral Researchers Program. H.W., S.O., and H.C.P. thank the Yukawa Institute for Theoretical Physics, Kyoto University, where we started the project during the YITP workshop YITP-T-24-03 on “Recent Developments and Challenges in Topological Phases.”
Appendix A Details of the Dirac trick
Appendix B Details of the honeycomb model
Tight-binding form—Consider a chiral symmetric honeycomb model of spinless fermions. Let the nearest-neighbor hopping be $t_{1}$. Assuming chiral symmetry represented by $\sigma_{3}$ in the sublattice space, the next allowed hopping $t_{2}$ is between third nearest neighbors. In momentum space, the Bloch Hamiltonian is given by $h_{\bm{k}}=Q_{\bm{k}}\sigma^{+}+Q_{\bm{k}}^{*}\sigma^{-}$, where
Quantum metric and density fluctuations energy—While the quantum metric is well-defined for isolated bands, it could become singular at band touching points. Indeed, in our F4 honeycomb model one finds $g_{ij}(\bm{q})\sim O(1/\delta q^{2})$ when $\bm{q}=\bm{q}_{\rm K}+\delta\bm{q}$ is around the K point (and similarly for the K’ point). More explicitly, we can parameterize $\bm{q}=\bm{q}_{K}+\delta q(\cos\theta,\sin\theta)$ and compute the leading singularity as
@@ Appendix B Details of the honeycomb model
We can now study how this singularity affects the SMA energy. For definiteness, we take $|\Phi\rangle$ to be the state obtained by filling all the negative-energy modes. As a technical trick, we further assume a finite linear size $L$ which is not divisible by $3$. This avoids sampling the singular K and K’ points in the finite-size momentum mesh and leads to a unique ground state. The local density operator is given by $\hat{n}_{\bm{R}}=\sum_{\sigma=A,B}\hat{c}_{{\bm{R}},\sigma}^{\dagger}\hat{c}_{{\bm{R}},\sigma}$. In the momentum-space, for $\bm{k}\neq\bm{0}$ and using $\hat{f}_{\bm{q},c}|\Phi\rangle=\hat{f}_{\bm{q},v}^{\dagger}|\Phi\rangle=0$, we have
@@
The leading term in the $\delta k\rightarrow 0$ limit of Eq. (16) is given by $t\sum_{\bm{q}}(2e_{\bm{q}})g_{ij}(\bm{q})\delta k_{i}\delta k_{j}$. Since $e_{\bm{q}}\sim O(\delta q^{2})$ around the K and K’ points, the divergence in the quantum metric is compensated and one finds $\langle\delta\bm{k}|\hat{H}|\delta\bm{k}\rangle\sim O(1)\delta k^{2}$, agreeing with general results derived form a double-commutator bound assuming only the locality and charge conservation symmetry of $\hat{H}$ [59, 60, 61, 62]. In typical SMA calculations, the normalization factor starts at $O(1)$ and so the SMA energy is gapless and disperses quadratically.
@@
$\begin{split}\int_{\frac{1}{L}}^{q_{0}}d\delta q\,\delta q\int_{0}^{2\pi}d\theta\,g_{\rm K}(\delta\bm{q})=\frac{2\pi}{\sqrt{3}}\log\left(q_{0}L\right)\left(\begin{array}[]{cc}1&\frac{1}{2}\\ \frac{1}{2}&1\end{array}\right),\end{split}$ (18)
@@
Therefore, although both the numerator and denominator in the SMA energy start with a quadratic term $\sim\delta k^{2}$, the coefficient for the numerator is $O(1)$ whereas that for the denominator is $O(\log L)$. This leads to $E^{\rm SMA}(\delta\bm{k})\sim O(1/\log L)$ in the limit $\delta k\rightarrow 0$, giving rise to an anomalous gap (Fig. 3(a)).
Interacting bound of the neutral excitation gap— We first Fourier transform the interaction $\hat{U}$ to find
@@
Next, we consider how the interaction modifies the charge-neutral excitation gap. While the particle-hole excitations $|\bm{k},\bm{q}\rangle$ are eigenstates of the free-fermion Hamiltonian, they become coupled when the interaction $\hat{U}$ is added to the Hamiltonian. As translation symmetry is preserved, sectors labeled by $\bm{k}$ remain decoupled. Let $\hat{H}^{\rm ph}|_{\bm{k}=\bm{0}}$ be the restriction of the Hamiltonian into the zero-momentum sector, defined by the matrix elements $\langle\bm{0},\bm{q}|\hat{H}^{\rm int}|\bm{0},\bm{q}^{\prime}\rangle$. Fig. 2(b) plots the low-lying eigenvalues of $\hat{H}^{\rm ph}|_{\bm{k}=\bm{0}}$ for different system sizes $L$. Fig. 3(b) shows the $O(1/L^{2})$ fit. We can now find an upper bound to the smallest eigenvalue through
[…]